# Energy-efficient recurrence quantification analysis

### v1

@@ 1 Introduction
The standard RQA algorithm begins with calculating the RP, which has a computational complexity of $\mathcal{O}(N^{2})$ Marwan et al. (2007). The lines formed by consecutive points in the RP are then detected, and their lengths are measured. As we consider diagonal and vertical lines, the computational complexity for each of these steps is also of order $\mathcal{O}(N^{2})$ (also in the case of symmetric RPs, where we would consider only one triangle of the RP). This complexity requires many calculation steps and, therefore, also increases calculation times. In the last decade, various approaches have been proposed to speed up RQA calculations. Classical approaches utilise parallelisation and use multiple graphics processing unit (GPU) devices Rawald et al. (2015, 2017). The latter distributes the calculations of RPs and line lengths on different GPU devices using the divide and recombine paradigm Rawald et al. (2014). Although the acceleration of the computations is remarkable, the complexity of the calculations has not changed. An approach that can indeed reduce the computational complexity extremely is a numerical and geometrical approximative ansatz, which can heavily accelerate the calculations by several magnitudes, but with the costs of increasing uncertainties Schultz et al. (2015); Spiegel et al. (2016). More recently, the introduction of microstate recurrence analysis Corso et al. (2018) has offered another (random sampling-based) approximative but fast approach, with much less uncertainty in the RQA measures da Cruz et al. (2025). Here, the line length distributions, required for the RQA measures, are estimated from randomly sampled sub-matrices from the RP.
@@
In the following, we will consider the original approach of RQA calculations, without parallelisation (although it would be possible) and without numerical/ geometrical approximation, but using the diagonal line structures of interest and developing some optimisation strategies. This optimisation follows the action numbers M2 (Reduce and compress data having the anticipated scientific value of the retained information and the resource requirements in mind) and M7 (“Design software for optimized energy consumption and provide tools to measure it”) Bruers et al. (2024). For the sake of simplicity, only the diagonal lines are considered here; the same approach can be applied for calculating the measures based on the vertical lines or even more complex patterns, such as triangle motifs (required for recurrence network measures Marwan et al. (2009)). The first optimisation will be very simple and has surely already been applied in several implementations. The second one is inspired by the microstates approach by da Cruz et al. (2025). These approaches are compared to a simple, straightforward standard implementation without any optimisations, as someone lacking knowledge of algorithmic and language-specific optimisations (such as those available in Julia) would perform it. The algorithms are implemented using the Julia language because it enables easy measurement of the performance of the implementations.
@@ 3 Optimisation Strategies
The following implementations have been tested on a single core of the high-performance cluster “Eunice Newton Foote” at the Potsdam Institute of Climate Impact Research (AMD EPYC 9554 processors with scalar frequencies of up to 3.75 GHz and 6 GByte DDR5 memory per core). The performance of an implementation was tested using the Julia macros @time and @timed. The calculations were performed 500 times to ensure reliable statistical estimates of the performance measurements.
@@ 3.1 Histogram Estimation without RP (RQA_woRP)
This kind of optimising the histogram calculation without calculating the RP beforehand is surely not novel. It has been used by the author before and surely might also have been implemented by others. For example, such implementation is available by Marwan (2025b) and was used in the run-time comparisons in the studies by Rawald et al. (2014, 2017) and Spiegel et al. (2016).
@@
Implementation Execution time Speedup Allocations Energy
RQA from RP (RQA_RP) $2.06\text{\,}\mathrm{s}$ 1 7 $28.6\text{\,}\mathrm{mWh}$
RQA without RP (RQA_woRP) $0.89\text{\,}\mathrm{s}$ 2.3 3 $12.4\text{\,}\mathrm{mWh}$
RQA without RP1 (RQA_woRP1) $0.48\text{\,}\mathrm{s}$ 4.3 3 $6.7\text{\,}\mathrm{mWh}$
RQA from sampling (RQA_Samp) $0.26\text{\,}\mathrm{s}$ 7.9 3 $3.6\text{\,}\mathrm{mWh}$
RQA from sampling2 (RQA_Samp2) $0.013\text{\,}\mathrm{s}$ 158 3 $0.18\text{\,}\mathrm{mWh}$
RQA from microstates $0.018\text{\,}\mathrm{s}$ 114 29 $0.25\text{\,}\mathrm{mWh}$
@@ 3.2 Sampling-based Histogram Estimation (RQA_Samp)
Based on the histogram estimation directly from the data, and without RP, we can further reduce the calculation costs by applying a random sampling scheme. Instead of sequentially testing every index $j$ from 1 to $N$ for every $i$, we can randomly draw indices $i$ and $j$ and test whether this pair fulfills the conditions given by Eqs. (3) and (4), i.e., correspond to the beginning of a diagonal line. If not, we randomly draw a new pair $i$ and $j$. If yes, we can measure the line length as is done in RQA_woRP, but stop after the end of a line and randomly sample a new line. By this sampling schema, we get only a subset of lines in the histogram $P(\ell)$, depending on the number $M$ of samplings. Using a smaller number of $M$, e.g., $M=4N$, we can significantly reduce the number of calculations to estimate $P(\ell)$. The resulting histogram $P(\ell)$ is a direct approximation of the histogram as it would be obtained using the RP_woRP algorithm described in Subsect. 3.1. Importantly, this approximation concerns the same underlying distribution, and the standard RQA measures can be computed from $P(\ell)$ without any modification. Even for small values of $M$, this distributional approximation is already quite accurate.
@@
Comparing to the standard estimation algorithm (RQA_RP) and the direct one (RP_woRP), we find a further speed-up depending on the number $M$ (Tab. 1). The memory allocations are the same as for the direct histogram estimation without RPs. A large number of samples $M=4N$, the calculation time is halved; further reducing $M$ to 20% of the data length speeds up the RQA estimation by almost two orders of magnitude.
Therefore, it might be desired to have a small $M$, but it obviously causes worse estimates of the line length measures. Considering the deviation $DET_{\text{sampled}}-DET_{\text{true}}$ shows that increasing $M$ reduces the error of the DET estimation when obtained with this sampling schema (Fig. 2). For the considered data with $N=25,000$, a sampling size of $M=500$ would already result in excellent estimations of DET with deviations of $10^{-4}$, using $M=5,000$ the deviations drop below $10^{-5}$.
Figure 2: Deviation values $DET_{\text{sampled}}-DET_{\text{true}}$ derived from sampled line length histogram for the Rössler system with 25,000 data points and for sampling size ranging from $M=10$ to $100,000$. The orange region indicates the deviation range of $\pm 10^{-4}$. A random sampling of only 500 line segments already results in minor deviations smaller than $10^{-4}$.
@@ 3.3 Efficiency by computation time
We also note that for small data ($N<10,000$), the sampling schema is not necessarily faster than RQA_woRP. Additionally, uncertainties in the RQA results increase for shorter time series when using sampling-based approaches, affecting both RQA_Samp and the microstates method (see appendix, Fig. 5), suggesting that RQA_woRP is preferable in such cases.
Using the same number $M$ in the recurrence microstates approach, the computation time is, in general, smaller (although it might not be directly comparable, because the microstates calculation is based on compiled C code highly optimised for Julia UFPR (2025)), but with the cost of less accurate results (see appendix, Figs. 4 and 5).
@@ 4 Discussion and conclusion
The choice of which approach to use depends on the research questions, data characteristics, and computational constraints. For small datasets ($N<10,000$) and when RPs are not needed, the RQA_woRP approach is the best choice, ensuring high accuracy and reasonable computation time. RQA_Samp becomes advantageous for longer time series, due to its linear sampling complexity ($M\propto N$), which scales more favorably than the $\mathcal{O}(N^{2})$ complexity of RQA_woRP, i.e., the computational savings grow with data length.
The results presented here demonstrate typical behaviour for a continuous chaotic system (Rössler attractor). Different dynamical regimes may affect the performance of these approaches differently. For instance, discrete maps like the logistic map generate RPs with shorter, more heterogeneous line structures. This leads to faster sampling (as valid line starts are found more frequently) but potentially larger uncertainties, particularly when $M$ is insufficient to adequately represent the line length distribution (Fig. 6A, C). In contrast, periodic systems have more homogeneous RPs with very long, non-interrupted, but less frequent diagonal lines. This reduces the probability of finding lines through random sampling, increasing computation time (can be even worse than the standard RQA_RP approach). However, the limited diversity of line lengths ensures accurate results even for small $M$ (Fig. 6B, D).
Sampling approaches are most efficient for homogeneously structured RPs characteristic of stationary dynamics. Non-stationary systems with heterogeneous RPs are expected to lead to larger uncertainties in RQA_Samp results, requiring either the use of RQA_woRP or increasing $M$. Systematic guidelines for selecting $M$ in non-stationary cases warrant future investigation.
@@ Acknowledgements
The author thanks Felipe Eduardo Lopes da Cruz, Thiago de Lima Prado, and Sergio Roberto Lopes for inspiring discussions on sampling microstates and for motivating this research. The author gratefully acknowledges support from the German Academic Exchange Service (DAAD) and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) via the project “Recurrence quantifiers as features for machine-learning-decision-making processes” (Grant No. 57705568), as well as from the Ministry of Research, Science and Culture (MWFK) of Land Brandenburg for supporting this project by providing resources on the high-performance computer system “Eunice Newton Foote” at the Potsdam Institute for Climate Impact Research (Grant No. 22-Z105-05/002/001).
@@
Figure 4: Deviation of DET derived from recurrence microstates approximation for the Rössler system with 25,000 data points and sampling size ranging from $M=10$ to $100,000$. The orange region indicates the deviation range of $\pm 10^{-4}$. To achieve deviations below $10^{-4}$, the sampling size must be at least $40,000$. Up to $M=5,000$, the estimated DET remains at $1.00$, resulting in the constant deviation.
Figure 5: Deviation $DET_{\text{sampled}}-DET_{\text{true}}$ for DET estimations using RQA_Samp, RQA_Samp2, and recurrence microstates approximation for the Rössler system for increasing data length NN. The sampling was $M=100,000$ for RQA_Samp and microstates RQA, and $M=5,000$ for RQA_Samp2. Even for a very small sampling number, the RQA_Samp approach has small deviations $10^{-4}$.
Figure 6: (A, C) Deviation $DET_{\text{sampled}}-DET_{\text{true}}$ and (B, D) computation times as in Figs. 2 and 3 for the (A, B) the logistic map and (C, D) a periodic signal. In (A, C) results for $N=25,000$ and using RQA_Samp2 are shown.

### v2

@@ 1 Introduction
The standard RQA algorithm begins with calculating the RP, which has a computational complexity of $\mathcal{O}(N^{2})$ Marwan et al. (2007). The lines formed by consecutive points in the RP are then detected, and their lengths are measured. As we consider diagonal and vertical lines, the computational complexity for each of these steps is also of order $\mathcal{O}(N^{2})$ (also in the case of symmetric RPs, where we would consider only one triangle of the RP). This complexity requires many calculation steps and, therefore, also increases computation times. In the last decade, various approaches have been proposed to speed up RQA calculations. Classical approaches utilise parallelisation and use multiple graphics processing unit (GPU) devices Rawald et al. (2015, 2017). The latter distributes the calculations of RPs and line lengths on different GPU devices using the divide and recombine paradigm Rawald et al. (2014). Although the acceleration of the computations is remarkable, the complexity of the calculations has not changed. An approach that can indeed reduce the computational complexity extremely is a numerical and geometrical approximative ansatz, which can heavily accelerate the calculations by several magnitudes, but with the costs of increasing uncertainties Schultz et al. (2015); Spiegel et al. (2016). More recently, the introduction of microstate recurrence analysis Corso et al. (2018) has offered another (random sampling-based) approximative but fast approach, with much less uncertainty in the RQA measures da Cruz et al. (2025); Ferreira et al. (2025). Here, the line length distributions, required for the RQA measures, are estimated from randomly sampled sub-matrices from the RP.
@@
In the following, we will consider the original approach of RQA calculations, without parallelisation (although it would be possible) and without numerical/ geometrical approximation, but using the diagonal line structures of interest and developing some optimisation strategies. This optimisation follows the action numbers M2 (“Reduce and compress data having the anticipated scientific value of the retained information and the resource requirements in mind”) and M7 (“Design software for optimized energy consumption and provide tools to measure it”) Bruers et al. (2024). These actions are part of the “Call-to-action in digital transformation”, which outline concrete measures developed by members of the ErUM-Data community during a dedicated workshop, aiming to reduce greenhouse gas emissions and promote the use of renewable energy within data-intensive research Bruers et al. (2024). For the sake of simplicity, only the diagonal lines are considered here; the same approach can be applied for calculating the measures based on the vertical lines or even more complex patterns, such as triangle motifs (corresponding to the recurrence configuration that two neighbours of a state vector are also neighbours, which is required for recurrence network measures Marwan et al. (2009); Donner et al. (2010)). The first optimisation will be very simple and has surely already been applied in several implementations. The second one is inspired by the microstates approach by da Cruz et al. (2025). These approaches are compared to a simple, straightforward standard implementation without any optimisations, as someone lacking knowledge of algorithmic and language-specific optimisations (such as those available in Julia) would perform it. The algorithms are implemented using the Julia language because it enables easy measurement of the performance of the implementations.
@@ 3 Optimisation Strategies
The following implementations have been tested on a single core of the high-performance cluster “Eunice Newton Foote” at the Potsdam Institute of Climate Impact Research (AMD EPYC 9554 processors with scalar frequencies of up to 3.75 GHz and 6 GByte DDR5 memory per core). The performance of an implementation was tested using the Julia macros @time and @timed. The calculations were performed 100 times to ensure reliable statistical estimates of the performance measurements.
@@ 3.1 Histogram Estimation without RP (RQA_woRP)
This kind of optimising the histogram calculation without calculating the RP beforehand is surely not novel. It has been used by the author before Marwan (2025b) and also have been implemented by others (e.g., Martinovič and Zitzlsberger (2019)). For example, such implementation is available in Marwan (2025b) and was used in the run-time comparisons in the studies by Rawald et al. (2014, 2017) and Spiegel et al. (2016).
@@
Implementation Computation time Speedup Allocations Energy
RQA from RP (RQA_RP) $2.01\text{\,}\mathrm{s}$ 1 7 $27.9\text{\,}\mathrm{mWh}$
RQA without RP (RQA_woRP) $0.88\text{\,}\mathrm{s}$ 2.3 3 $12.2\text{\,}\mathrm{mWh}$
RQA without RP1 (RQA_woRP1) $0.48\text{\,}\mathrm{s}$ 4.2 3 $6.7\text{\,}\mathrm{mWh}$
RQA from sampling (RQA_Samp) $0.25\text{\,}\mathrm{s}$ 8.0 3 $3.5\text{\,}\mathrm{mWh}$
RQA from sampling2 (RQA_Samp2) $0.012\text{\,}\mathrm{s}$ 168 3 $0.17\text{\,}\mathrm{mWh}$
RQA from microstates $0.017\text{\,}\mathrm{s}$ 1148 29 $0.24\text{\,}\mathrm{mWh}$
@@ 3.2 Sampling-based Histogram Estimation (RQA_Samp)
Based on the histogram estimation directly from the data, and without RP, we can further reduce the calculation costs by applying a random sampling scheme. Instead of sequentially testing every index $j$ from 1 to $N$ for every $i$, we can randomly draw indices $i$ and $j$ and test whether this pair fulfils the conditions given by Eqs. (3) and (4), i.e., correspond to the beginning of a diagonal line. If not, we randomly draw a new pair $i$ and $j$. If yes, we can measure the line length as is done in RQA_woRP, but stop after the end of a line and randomly sample a new line. By this sampling schema, we get only a subset of lines in the histogram $P(\ell)$, depending on the number $M$ of samplings. Using a smaller number of $M$, e.g., $M=4N$, we can significantly reduce the number of calculations to estimate $P(\ell)$. The resulting histogram $P(\ell)$ is a direct approximation of the histogram as it would be obtained using the RP_woRP algorithm described in Subsect. 3.1. Importantly, this approximation concerns the same underlying distribution, and the standard RQA measures can be computed from $P(\ell)$ without any modification. Even for small values of $M$, this distributional approximation is already quite accurate.
@@
Comparing to the standard estimation algorithm (RQA_RP) and the direct one (RP_woRP), we find a further speed-up depending on the number $M$ (Tab. 1). The memory allocations are the same as for the direct histogram estimation without RPs. A large number of samples $M=4N$, the computation time is halved; further reducing $M$ to 20% of the data length speeds up the RQA estimation by two orders of magnitude.
Therefore, it might be desired to have a small $M$, but it obviously causes worse estimates of the line length measures. Considering the error $DET_{\text{sampled}}-DET_{\text{true}}$ shows that increasing $M$ reduces the error of the DET estimation when obtained with this sampling schema (Fig. 2). For the considered data with $N=25,000$, a sampling size of $M=500$ would already result in excellent estimations of DET with errors of $10^{-4}$, using $M=5,000$ the estimation errors drop below $10^{-5}$.
Figure 2: Estimation error $DET_{\text{sampled}}-DET_{\text{true}}$ derived from sampled line length histogram for the Rössler system with 25,000 data points and for sampling size ranging from $M=10$ to $100,000$. The orange region indicates the error range of $\pm 10^{-4}$. A random sampling of only 500 line segments already results in minor errors smaller than $10^{-4}$.
@@ 3.3 Efficiency by computation time
We also note that for small data ($N<10,000$), the sampling schema is not necessarily faster than RQA_woRP. Additionally, uncertainties in the RQA results increase for shorter time series when using sampling-based approaches, affecting both RQA_Samp and the microstates method (see appendix, Fig. A2), suggesting that RQA_woRP is preferable in such cases.
Using the same number $M$ in the recurrence microstates approach Ferreira et al. (2025), the computation time is, in general, smaller (although it might not be directly comparable, because the microstates calculation is based on compiled C code highly optimised for Julia UFPR (2025)), but with the cost of less accurate results (see appendix, Figs. A1 and A2).
@@ 4 Discussion and conclusion
The choice of which approach to use depends on the research questions, data characteristics, and computational constraints. For small datasets ($N<10,000$) and when RPs are not needed, the RQA_woRP approach is the best choice, ensuring high accuracy and reasonable computation time. RQA_Samp becomes advantageous for longer time series, due to its linear sampling complexity ($M\propto N$), which scales more favourably than the $\mathcal{O}(N^{2})$ complexity of RQA_woRP, i.e., the computational savings grow with data length.
The results presented here demonstrate typical behaviour for a continuous chaotic system (Rössler attractor). Different dynamical regimes may affect the performance of these approaches differently. For instance, discrete maps like the logistic map generate RPs with shorter, more heterogeneous line structures. This leads to faster sampling (as valid line starts are found more frequently) but potentially larger uncertainties, particularly when $M$ is insufficient to adequately represent the line length distribution (Fig. A3A, C). In contrast, periodic systems have more homogeneous RPs with very long, non-interrupted, but less frequent diagonal lines. This reduces the probability of finding lines through random sampling, increasing computation time (can be even worse than the standard RQA_RP approach). However, the limited diversity of line lengths ensures accurate results even for small $M$ (Fig. A3B, D).
Sampling approaches are most efficient for homogeneously structured RPs, characteristic of stationary dynamics. Non-stationary systems with heterogeneous RPs are expected to lead to larger uncertainties in RQA_Samp results, requiring either the use of RQA_woRP or increasing $M$. Systematic guidelines for selecting $M$ in non-stationary cases warrant future investigation.
@@ Acknowledgements
The author thanks Felipe Eduardo Lopes da Cruz, Thiago de Lima Prado, and Sergio Roberto Lopes for inspiring discussions on sampling microstates and for motivating this research. The author gratefully acknowledges support from the German Academic Exchange Service (DAAD) and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) via the project “Recurrence quantifiers as features for machine-learning-decision-making processes” (Grant No. 57705568), as well as from the Ministry of Research, Science and Culture (MWFK) of Land Brandenburg for supporting this project by providing resources on the high-performance computer system “Eunice Newton Foote” at the Potsdam Institute for Climate Impact Research (Grant No. 22-Z105-05/002/001). For the RQA estimations using the microstates approach, the recently published Julia package RecurrenceMicrostatesAnalysis.jl Ferreira et al. (2025) has been used.
@@
Performance of the algorithm and further dynamical systems
Figure A1: Estimation error of DET derived from recurrence microstates approximation for the Rössler system with 25,000 data points and sampling size ranging from $M=10$ to $100,000$. The orange region indicates the error range of $\pm 10^{-4}$. To achieve estimation errors below $10^{-4}$, the sampling size must be at least $40,000$. Up to $M=5,000$, the estimated DET remains at $1.00$, resulting in the constant deviation.
Figure A2: Estimation error $DET_{\text{sampled}}-DET_{\text{true}}$ for DET estimations using RQA_Samp, RQA_Samp2, and recurrence microstates approximation for the Rössler system for increasing data length NN. The sampling was $M=100,000$ for RQA_Samp and microstates RQA, and $M=5,000$ for RQA_Samp2. Even for a very small sampling number, the RQA_Samp approach has small estimation errors $10^{-4}$.
Figure A3: (A, C) Estimation error $DET_{\text{sampled}}-DET_{\text{true}}$ and (B, D) computation times as in Figs. 2 and 3 for the (A, B) the logistic map and (C, D) a periodic signal. In (A, C) results for $N=25,000$ and using RQA_Samp2 are shown.
Recommendation for required data length
While a more systematic study is necessary to get good recommendations on required time series lengths for the proposed RQA_samp2 approach, we can already gather some clues.
Errors in estimating the RQA measure and computation times with RQA_samp2 depend on the characteristics of the data. Data that cause relatively homogeneously distributed long lines in an RP, such as for Rössler or Lorenz systems, show estimation errors that are already quite small for short data lengths, with $100<N<1,000$. The speedup in computation time becomes significant for $N>1,000$ (Fig. A4A). For systems that generate many short lines and single points, as typical for maps, the errors are generally larger, but the computation time benefits already significantly for small $N$ (Fig. A4B). To achieve reasonable results, at least $N>2,000$ are necessary, although this number can be lowered by increasing the sampling size $M$. In contrast, sparse RPs with consistent patterns cause numerous unsuccessful sampling tries, leading to limited speedup but very slight errors (Fig. A4C). Therefore, this approach might be less effective for periodic systems compared to signals with more complex or irregular dynamics.
Figure A4: Speedup of computation time ($t_{\text{RQA\_samp}^{2}}/t_{\text{RQA\_woRP}}$) and absolute estimation error $|DET_{\text{sampled}}-DET_{\text{true}}|$ as function of data length using the RQA_samp2 approach for (A) the Rössler system, (B) the logistic map, and (C) a periodic signal. Shaded areas indicate the 5% and 95% quantiles obtained from 500 repetitions; solid lines denote the median.